# A Method for On‑Orbit Calibration of the VLAST‑P Electromagnetic Calorimeter

### v2

@@ 4.2 Event Selection
Due to arbitrary incident direction of cosmic ray particles, it is necessary to distinguish the events that can be used for calibration. First, according to the trigger logic, during the MIP calibration process, it is required that the hit signals from multiple detectors be combined using a logical AND operation. This ensures that only events where several detectors are simultaneously triggered are selected for calibration. A more specific set of thresholds is applied in the simulation, requiring a logical AND signal between the two ACD layers, as well as valid energy depositions in both the converter region of the tracker and the calorimeter. An event is kept if at least two units in two ACD layers, one for each, have more than 0.6 MeV of deposited energy, one converter unit has more than 2.52 MeV, and one ECAL cell has more than 36 MeV. These thresholds help to ensure that the signals come from real particle interactions rather than random background. The input spectrum used to simulate the realistic spatial cosmic ray environment is generated from the database described earlier. From these input spectra, a pronounced east–west effect is observed, as shown in figure 4.
@@
For particles penetrating the detector at large incident angles, the path length calculation can be complicated by shower development. To determine the optimal selection criterion, we analyzed the hit cell multiplicity distributions for both showering and non-showering events. Based on this study, a threshold of fewer than 6 hit cells was established. This cut effectively discriminates against showering events, ensuring that the selected tracks have well-defined path lengths within the crystals. Furthermore, a correction for the path length has been applied in this step by evaluating the true trajectory using the incident angle and initial position of the particle. There lies a tracker in front of the ECAL. In the real on-orbit calibration, with the reconstruction algorithm of the tracker, the incident angle can also be used to select events. Additionally, in some events, it is inevitable that some hits come from a very short path length. To exclude these, hits with a path length less than 2 cm are not counted.
@@ 4.3 Performance
The selection criteria are primarily designed to reject events caused by particles entering at large incident angles, thereby retaining events suitable for calibration purposes. As presented in table 1, each selection stage leads to a substantial reduction in the number of surviving events. The table also lists the selection efficiency at each stage and the cumulative efficiency up to that point. Starting from the full simulated dataset, the event count decreases step-by-step, ultimately retaining only 2.95% of the total events with a source filter mentioned in Section 4.1. This result demonstrates that the selection strategy is effective in suppressing showering events while preserving a clean sample of physical events suitable for further analysis.
@@
On orbit, the satellite can operate in a sun-pointing mode, in which the front side of the detector continuously faces the Sun. Under this configuration, the detector alternately points toward the west and east directions with an approximately equal probability. Since the incident particle rigidity spectrum can differ significantly between the two orientations, the rigidity spectra shown in figure 4 are separately adopted as input distributions in the simulation to evaluate their impact. The analysis indicates that the MPV of the energy deposition for protons arriving from the East and West directions exhibits a discrepancy of approximately 6%, as illustrated in figure 5.
@@
A path length correction is applied in the analysis to compensate for the effect of particle incident angles on the measured energy deposition. Particles entering the detector at oblique angles traverse a longer path through the scintillator material and may penetrate multiple crystal bars, resulting in larger apparent energy depositions compared with perpendicular incidence. Without this correction, the reconstructed energy for large-angle events would be systematically underestimated due to the significant number of oblique incidences. After applying the path length correction, the energy spectrum exhibits a Landau-like shape, and the MIP peak position becomes more consistent across different angular ranges, improving both the accuracy and stability of the calibration.
@@
In the actual space environment, secondary particles, most of which are protons, also exist under space conditions, typically with energies below $10\,\mathrm{GeV}$. These secondary particles originate from interactions between high-energy primary cosmic rays and atmospheric atoms, producing cascades of new particles through nuclear reactions. According to the AMS proton flux measurements, the integral flux of secondary protons accounts for approximately 40% of the total integral flux of all protons in the equatorial region [SecondaryProton]. In this study, secondary protons are modeled using the kinetic energy distribution measured by AMS, with arrival directions sampled isotropically from a source sphere. The simulation results show that the hit multiplicity distribution follows an approximately exponential decay, indicating that most secondary protons deposit energy in only a few crystals. The energy distribution of secondary protons is broader compared to that of primary protons. The selection efficiency of secondary protons is about half of the primary efficiency. These secondary events are combined with the primary proton events in proportion to their respective integral fluxes.
The on-orbit MIP calibration results show that the energy deposition spectrum of protons follows a Landau-like distribution, with the most probable value (MPV) measured at $112.9\,\mathrm{MeV}$ in a single CsI(Tl) unit, as shown in figure 6(a). In the case of helium nuclei (He), it should be emphasized that their flux in space is also considerable, making them suitable for detector calibration. Compared with protons, helium nuclei are almost free from secondary components, as the production of secondary helium through nuclear interactions is extremely rare. A comparable distribution is obtained for helium nuclei, whose MPV is $443.4\,\mathrm{MeV}$, nearly four times that of the proton MIP, depicted in figure 6(b). This scaling behavior can be naturally understood from the Bethe–Bloch formalism, where the stopping power of charged particles in matter is approximately proportional to the square of their charge number ($Z^{2}$). As helium nuclei carry twice the charge of protons ($Z=2$), their energy deposition is therefore expected to be enhanced by a factor of about four, in good agreement with the observed ratio. Furthermore, observations suggest that secondary cosmic rays mainly come from regions near the sources and from the general interstellar medium [SecondaryGeneral].
@@ 5 Conclusion And Outlook
To quantify the on-orbit calibration time requirement, the incident proton flux was taken as $\Phi_{p}=0.015\ \mathrm{cm^{-2}\ s^{-1}\ sr^{-1}}$, representative of the near-equatorial geomagnetic cutoff conditions. For each orbital revolution, an effective observation window of approximately $t_{\mathrm{eff}}=1200\ \mathrm{s}$ is available within the target geomagnetic latitude range ($20^{\circ}<\mathrm{lat}<20^{\circ}$) [DAMPE20], based on an orbital period of $T_{\mathrm{orbit}}=1.5\ \mathrm{h}=5400\ \mathrm{s}$. Based on Monte Carlo simulations employing a spherical source with a radius of $0.5\,\mathrm{m}$, the geometric factor of the electromagnetic calorimeter for MIP calibration is calculated to be $G=6.34\times 10^{-3}\penalty 10000\ \mathrm{m^{2}\cdot sr}$. Taking into account the proton flux and the orbital period mentioned above, this results in an expected yield of approximately 1142 valid calibration events per orbit. Considering the combined flux of primary and secondary protons, a sample of approximately $3\times 10^{3}$ events per channel is required across the 25 readout channels to constrain the statistical uncertainty to $0.5\%$. In total, approximately $7.5\times 10^{4}$ events are required. Consequently, the total acquisition requirement amounts to approximately 65.7 orbital periods, corresponding to about 98 hours ($\approx 4.1$ days) of effective data-taking. This estimate serves as a practical baseline for scheduling and planning the calibration campaigns during the operational phase of the mission.

### v3

@@ 4.2 Event Selection
Calibration events are selected using a strict four-fold coincidence logic requiring valid hits in both ACD layers, the converter, and the ECAL. The specific energy thresholds are set to $0.6\,\mathrm{MeV}$ for each ACD layer, $2.52\,\mathrm{MeV}$ for single converter tile, and $36\,\mathrm{MeV}$ for single ECAL crystal, ensuring the effective rejection of background noise. The input spectrum used to simulate the realistic spatial cosmic ray environment is generated from the database described earlier. From these input spectra, a pronounced east–west effect is observed for positive particles, as shown in figure 5.
@@
For particles penetrating the detector at large incident angles, the path length calculation can be complicated by shower development. To determine the optimal selection criterion, we analyzed the hit cell multiplicity distributions for both showering and non-showering events. Based on this study, a threshold of fewer than 6 hit cells was established. This cut effectively discriminates against showering events, ensuring that the selected tracks have well-defined path lengths within the crystals. Furthermore, a correction for the path length has been applied in this step by evaluating the true trajectory using the incident angle and initial position of the particle, as measured by the tracker located upstream of the ECAL. In the real on-orbit calibration, with the reconstruction algorithm of the tracker, the incident angle can also be used to select events. Additionally, in some events, it is inevitable that some hits come from a very short path length. To exclude these, hits with a path length less than 2 cm are not counted.
@@ 4.3 Performance
The selection criteria are primarily designed to reject events caused by particles entering at large incident angles, thereby retaining events suitablefor calibration purposes. As presented in Table 1, each selection stage leads to a substantial reduction in the number of surviving events. The table lists both the selection efficiency at each step and the cumulative efficiency. Starting from the full simulated dataset, the event count decreases sequentially as increasingly restrictive criteria are applied. Logic trigger selects events satisfying the detector’s trigger conditions. ECAL hit number is applied to suppress cluster shower events. Incident direction requires that the particle trajectory reconstructed by the tracker in the simulation is within 20° of the z-axis, ensuring nearly normal incidence. After the complete set of source-filter selections described in Section 4.1, only 2.95% of the total simulated events remain. This result demonstrates that the selection strategy is effective in suppressing showering events while preserving a clean sample of physical events suitable for further analysis.
@@
On orbit, the satellite can operate in a sun-pointing mode, in which the front side of the detector continuously faces the Sun. Under this configuration, the detector alternately points toward the west and east directions with an approximately equal probability. Since the incident particle rigidity spectrum can differ significantly between the two orientations, the rigidity spectra shown in figure 5 are separately adopted as input distributions in the simulation to evaluate their impact. The analysis indicates that the MPV of the energy deposition for protons arriving from the East and West directions exhibits a shift of approximately 6%, as illustrated in figure 6.
@@
A path length correction is applied in the analysis to compensate for the effect of particle incident angles on the measured energy deposition. Particles entering the detector at oblique angles traverse a longer path through the scintillator material and may penetrate multiple crystal bars, resulting in larger apparent energy depositions compared with perpendicular incidence. After applying the path length correction, the energy spectrum exhibits a Landau-like shape, and the MIP peak position becomes more consistent across different angular ranges, improving both the accuracy and stability of the calibration.
@@
In the actual space environment, secondary particles, most of which are protons, also exist under space conditions, typically with energies below $10\,\mathrm{GeV}$. These secondary particles originate from interactions between high-energy primary cosmic rays and atmospheric atoms, producing cascades of new particles through nuclear reactions. According to the AMS proton flux measurements, the integral flux of secondary protons accounts for approximately 40% of the total integral flux of all protons in the equatorial region [2]. In this study, secondary protons are modeled using the kinetic energy distribution measured by AMS, with arrival directions sampled isotropically from a source sphere. The simulation results show that the hit multiplicity distribution follows an approximately exponential decay, indicating that most secondary protons deposit energy in only a few crystals. The energy distribution of secondary protons is broader compared to that of primary protons. The selection efficiency of secondary protons is about half of the primary efficiency. These secondary events are combined with the primary proton events in proportion to their respective integral fluxes.
The on-orbit MIP calibration results show that the energy deposition spectrum of protons follows a Landau-like distribution, with the most probable value (MPV) measured at $112.9\,\mathrm{MeV}$ in a single CsI(Tl) unit, as shown in figure 7(a). In the case of helium nuclei (He), it should be emphasized that their flux in space is also considerable, making them suitable for detector calibration. Compared with protons, helium nuclei are almost free from secondary components, as the production of secondary helium through nuclear interactions is extremely rare. A comparable distribution is obtained for helium nuclei, whose MPV is $443.4\,\mathrm{MeV}$, nearly four times that of the proton MIP, depicted in figure 7(b). This scaling behavior can be naturally understood from the Bethe–Bloch formalism, where the stopping power of charged particles in matter is approximately proportional to the square of their charge number ($Z^{2}$). As helium nuclei carry twice the charge of protons ($Z=2$), their energy deposition is therefore expected to be enhanced by a factor of about four, in good agreement with the observed ratio.
@@ 5 Conclusion And Outlook
To quantify the on-orbit calibration time requirement, the incident proton flux was taken as $\Phi_{p}=0.015\ \mathrm{cm^{-2}\ s^{-1}\ sr^{-1}}$, obtained by approximately integrating the AMS-measured differential proton flux in the equatorial region [2]. This value is representative of near-equatorial geomagnetic cutoff conditions. For each orbital revolution, an effective observation window of approximately $t_{\mathrm{eff}}=1200\ \mathrm{s}$ is available within the target geomagnetic latitude range ($20^{\circ}<\mathrm{lat}<20^{\circ}$) [4], based on an orbital period of $T_{\mathrm{orbit}}=1.5\ \mathrm{h}=5400\ \mathrm{s}$. Based on Monte Carlo simulations employing a spherical source with a radius of $0.5\,\mathrm{m}$, the geometric factor of the electromagnetic calorimeter for MIP calibration is calculated to be $G=6.34\times 10^{-3}\penalty 10000\ \mathrm{m^{2}\cdot sr}$. Taking into account the proton flux and the orbital period mentioned above, this results in an expected yield of approximately 1142 valid calibration events per orbit. Considering the combined flux of primary and secondary protons, a sample of approximately $3\times 10^{3}$ events per channel is required across the 25 readout channels to constrain the statistical uncertainty to $0.5\%$. In total, approximately $7.5\times 10^{4}$ events are required. Consequently, the total acquisition requirement amounts to approximately 65.7 orbital periods, corresponding to about 98 hours ($\approx 4.1$ days) of effective data-taking. This estimate serves as a practical baseline for scheduling and planning the calibration campaigns during the operational phase of the mission.The dominant sources of systematic uncertainty are expected to arise from the on-orbit operational environment. For example, temperature variations may affect the detector response; according to communication with the satellite system team, the temperature variation within one orbit is controlled within 0.5 degree, which may introduce an uncertainty of about $1\%$ in the calibration precision. In addition, uncertainties in the geomagnetic field model and the effect of solar modulation can also contribute to systematic deviations. Based on the experience of the DAMPE Collaboration, these effects may introduce an additional uncertainty at the level of approximately $1\%$ [18].